Shift Summary for day, 3 Sept, 1999

Angela Bellavance (bellavan@rice.edu)
Fri, 3 Sep 1999 15:34:10 -0500 (CDT)

Shift Summary for day shift, 03/Sept/99

On Shift: A. Tripathi, A. Bellavance

Hours of beam delivered at all intensities: 5.3
Hours of beam delivered at nominal intensity: 5.3
Hours of E832 data (special runs not included): 4.3
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
ktev3 SCSI bus for tape drives hung
CsI channel 226 gain drift

========================================================================
Runs ended this shift:
Run 14420 - ? events : BEAM_832 (run aborted - no final info)
Run 14421-23 : unsuccessful
Run 14424 - 13.5M events : BEAM_832
Run 14425 - ongoing

Tapes brought to FCC this shift: none (Viv wants to take them in next
shift)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E832 datataking and tape labelling

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14420 198 7:03 Take over, getting beam.
7:12 Tape mount request: QKN033 in ktev4, tps3d2.
Tape already in drive, so just clicked OK.
Succeeded (I think).
7:20 Ed drops in (good thing he did!)
7:35 Tape mount request: QKA023 in ktev3, tps2d3.
Tape already in drive, so just clicked OK.
Failed.
225 7:38 DAQDEAD_SEC alarm. We missed part of spill 223.
227 7:43 Run abort!! Error from ktev3_drv2: "no provider
in service group." Tapes unloaded fine. Cleaned
drive. More problems. Eventually diagnosed as
SCSI bus for ktev3 drives is hung. Power cycled
crate for ktev3 tape drives.
8:20 Tape mount request: QKA023 in ktev3, tps2d3 again.
Decide to wait until ktev3 tape drives are working
again to click "ok".
8:26 TeV abort @ D1.
8:28 TeV abort @ D1.
8:51 SY energy conservation in effect.
14424 1 8:57 Start new run after recovering from ktev3 tape
problems. 3X kumquats hang. KFREND fixes.
8:59 Clicked "ok" for QKA023 tape mount. Success this
time (I think).
8:59 Message from Ch.13: Quench at A1. Est. 40 min.
14424 28 9:30 Beam returns! Position is off a little and
intensity is high for a few spills but settles out
on its own.
14424 37 9:45 Trigger hangs. Takes several KFRENDs to fix.
-40
14424 41 9:50 We're back, but SEM is high (7.1e12).
42 9:52 SEM is high (7.2e12). Call MCR. They will "work
on it."
64 10:21 Ktevana errors - KFREND.
10:27 Notice BOA has been high. Call MCR to complain.
90 10:55 Abort @ F1. Beam goes away.
10:58 Message from 13: C4 wet engine adjustment. Est.
10 min.
104 11:14 Beam returns!
11:22 BOA still bad. Call MCR. Again.
114-5 11:28 SY energy conserv.
121-3 11:37 SY energy conserv.
128 11:45 Ktevana errors - KFREND.
136 11:56 SY energy conserv.
140 12:04 SY energy conserv.
141 Weird beam structure. Looks like the beam stop is
in. Call MCR, they say they'll look at it but
never call us back. It corrects itself the next
spill anyway.
162 12:32 Abort.
13:26 With all requests to improve BOA from the shift
crew having failed, Ed calls MCR.
223 13:54 Stopped triggers to investigate ch. 226 gain drift
problem. Up by 11%.
225 We're alive again.
14424 251 14:31 Ed decides to stop run, start new with channel 226
corrected.
14425 1 14:38 Start new run.
6 14:46 Beam disappears. Message from Ch.13: "Problems
with supply V701. Est. 25 minutes."
15:18 Angela finishes labeling 240 tapes.
15:30 Hand off to eve shift.

========================================================================
Follow up Notes on Previous Shift

Miguel sent out the shift summary for yesterday's owl shift this morning,
but labeled it "eve".

========================================================================
Items to Follow up on Next Shift

Keep taking standard high-intensity 832 data.

For Arun, Angela

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke